# Benchmarking Fake Voice Detection in the Fake Voice Generation Arms Race

### v1

@@ Abstract
As advances in synthetic voice generation accelerate, an increasing variety of fake voice generators have emerged, producing audio that is often indistinguishable from real human speech. This evolution poses new and serious threats across sectors where audio recordings serve as critical evidence. Although fake voice detectors are also advancing, the arms race between fake voice generation and detection has become more intense and complex. In this work, we present the first large-scale, cross-domain evaluation of fake voice detectors, benchmarking 8 state-of-the-art models against datasets synthesized by 20 different fake voice generation systems. To the best of our knowledge, this is the most comprehensive cross-domain assessment conducted to date. Our study reveals substantial security vulnerabilities in current fake voice detection systems, underscoring critical gaps in their real-world robustness. To advance the field, we propose a unified and effective metric that consolidates the diverse and often inconsistent evaluation criteria previously used across different studies. This metric enables standardized, straightforward comparisons of the robustness of fake voice detectors. We conclude by offering actionable recommendations for building more resilient fake voice detection technologies, with the broader goal of reinforcing the foundations of AI security and trustworthiness.
@@ 1 Introduction
The evolution of fake voice generation has changed digital communication, allowing applications such as text-to-speech (TTS), voice conversion, and audio reconstruction to achieve unprecedented naturalness and quality [31]. Early techniques based on parametric [133, 89, 117, 118] or concatenative [77, 53, 111] methods have given way to deep neural architectures such as WaveNet [122] which introduced an autoregressive generative model for raw audio, and Tacotron [107] which employs sequence-to-sequence architectures to map text directly to spectrograms. Recently, vocoders based on generative adversarial networks (GAN) [33] and diffusion have pushed the boundaries further, achieving near-human levels of prosody and intelligibility [147, 57, 59].
This rapid evolution in synthesis quality has been applied a wide range of services, from virtual assistants and accessibility tools to customer service and content creation [109, 80]. However, as fake voice approaches human parity, it also exposes vulnerabilities in automatic speaker verification (ASV) systems and their related countermeasure (CM) systems [38, 144]. Fake voice generation poses a critical security risk in sensitive domains such as financial services, law enforcement, and biometric security [56, 50, 8]. Recent incidents have demonstrated how malicious actors can exploit high-fidelity voice conversion to impersonate executives and facilitate unauthorized transfers [1, 144].
@@
Consequently, the arm race between advanced fake voice generation and fake voice detection has emerged [19]. Although prior studies and benchmarks have evaluated isolated components [2, 147, 116, 84, 68], few have addressed the coupled evolution of fake voice generation methods and detection systems. Prior research in fake voice detection has conventionally aggregated fake voice samples into a single homogeneous dataset for detector training and evaluation [125, 73, 141, 94, 76, 129, 47, 128, 127]. For instance, in [140], they treat detection primarily as a static evaluation task, using a fixed dataset (constructed from fake voice samples) to assess the performance of various fake voice speech detectors across standardized manipulations and scenarios. This practice ignores critical nuances: it prevents the identification of unique acoustic artifacts inherent to each generation paradigm and masks variations in detector performance across generator types. Recent studies show that acoustic models and vocoders leave unique fingerprints in fake voice waveforms [21]. These fingerprints can be used for source attribution and to improve synthesis quality.
@@
We introduce an ecosystem-level benchmark that systematically evaluates the interplay between modern fake voice generation “engines” and detection “guardians.” Our framework challenges 20 state-of-the-art (SOTA) fake voice generation systems, including TTS, TTS with zero-shot voice conversion, and vocoder-based audio reconstructions, with 8 leading fake voice detectors using a novel one-to-one evaluation protocol. Unlike conventional assessments based on aggregated synthetic data, our approach evaluates each generator-detector pair individually to expose unique acoustic artifacts, method-specific sensitivities, and previously hidden vulnerabilities. Furthermore, by integrating explainability analysis, we clarify the underlying reasons for detector performance discrepancies, enabling the community to pinpoint the generation techniques that most threaten current defenses. Figure 1 outlines our SoK pipeline.
@@
    Comprehensive Overview: We provide a comprehensive taxonomy of fake voice generation and fake voice detection systems, framed through a security lens. We identify critical vulnerabilities in fake voice detection systems that failing against high-fidelity fake voices. We also pinpoint the lack of explainable analysis of the reasons behind the detection behaviors, highlighting urgent security gaps.
@@
    Refine Taxonomy of Fake Voice Detection: We refined the taxonomy structure by providing generalization-optimized models in addition to the regular taxonomy of the detector pipeline, pinpointing the generalization needs, which is also demonstrating in our cross-domain evaluation.
@@
    Empirical Security Assessment and its Rationale: We conduct a large-scale, cross-domain evaluation analysis involving 20 fake voices and 8 fake voice detectors. This study not only highlights significant security vulnerabilities and quantifies performance variations across diverse generator-detector interactions, but also explains why those interaction might happen. We provide critical insights into the root causes of these vulnerabilities. Such understanding is essential for developing more robust and transparent fake voice ecosystems.
@@
    Recommended Practices for Secure Systems: We propose concrete practices for fake voice detectors and also provide guides for the development of fake voice generations that balance high fidelity with security resilience.
[…]
Paper Year Type Modeling Params Open Control Dataset Inference Zero
(M) Src Prosody Style Spk Hours Languages Speed Shot
Tacotron 2017 AM AR $\times$ $\times$ $\times$
DeepVoice 2017 AM AR $\times$ $\times$ $\times$
TransformerTTS 2018 AM Trans $\times$ $\times$ $\times$
FastSpeech 2019 AM Trans 30.1 $\times$ $\times$
Grad-TTS 2021 AM Diff 7.6 $\times$ $\times$ $\times$
VITS 2021 Fully E2E VAE $\times$
Your-TTS 2021 Fully E2E Trans+VAE $\times$
NaturalSpeech 2022 Fully E2E VAE 28.7 $\times$ $\times$ $\times$ $\times$
NaturalSpeech2 2023 NC-LM Diff 400 $\times$
VALL-E 2023 NC-LM Trans $\times$
VoiceBox 2023 AM FM 330 $\times$
NaturalSpeech3 2024 NC-LM Trans+Diff 1000 $\times$
MaskGCT-TTS 2024 NC-LM Trans+VQ-VAE 695
@@ 2 Fake Voice Generators
In this section, we provide a taxonomy of fake voice generation systems, with a specific focus on Text-to-Speech (TTS). TTS refers to the process of converting written text into natural-sounding spoken audio. TTS with zero-shot voice conversion is a technique that enables voice generation in a target speaker’s voice using only a short reference audio clip, without requiring prior training on that specific voice [15]. There are several categories of fake voices. A high-level pipeline is shown in Figure 2.
@@
Figure 2: Fake Voice Generation Evolution. Traditional neural TTS pipelines convert text to acoustic representations, then utilize vocoders to synthesize waveforms. Fully E2E models follow a traditional approach, but optimize all components at once. Neural codec models encode and decode audio using a data-driven approach instead of static compression algorithms. Audio inputs are optional and are often used for speaker conditioning in zero-shot TTS with voice conversion.
From Figure 2, we can see two methods for arriving at the output fake voice. First, traditional approaches use a neural network-based text encoder to generate latent features based on input text. Then, a vocoder synthesizes the waveform from hidden features. Note that traditional approaches can also be end-to-end (E2E), where all components are optimized at once. Second, neural audio codecs provide a machine learning-based compression mechanism for raw audio. These approaches utilize the encoder and decoder from a neural audio codec, along with a text encoder, to synthesize raw audio. The rest of this section will expand on different methods of fake voice generation.
2.1 Traditional Generators with Vocoder
Traditional fake voice generation systems typically include two main parts: a text encoder that converts text inputs into meaningful acoustic features (the acoustic model), and a vocoder which synthesizes raw waveforms using the acoustic features.
2.1.1 Acoustic Model
Acoustic models generate acoustic representations from linguistic features or even directly from phonemes or characters [116]. Table 1 shows some of the acoustic models that we investigated, which span three main categories (more details in Table 9 of Appendix): autoregressive, flow matching & diffusion-based, and transformer-based.
Autoregressive Model: This type of acoustic model transformed fake voice generation by aggregating text-to-speech pipelines into neural networks and obtaining excellent naturalness with minimum manual feature engineering. Inspired by the Tacotron [102, 107, 134] series, these models autoregressively produce spectrograms from text using encoder-attention-decoder architectures. Later improvements included DurIAN [143] for stable alignment use duration predictors rather than attention. DeepVoice 1-3 [6, 85, 5] combined CNNs for mel spectrogram generation and linguistic feature extraction. Autoregressive systems remain fundamental despite difficulties in latency and controllability, usually coupled with non-autoregressive components such as duration predictors or equipped with advanced vocoders to balance quality and speed in contemporary systems.
Flow-Matching & Diffusion-Based Models: Flow-matching and diffusion models represent two powerful paradigms for transforming simple base distributions (e.g., Gaussians) into complex data distributions, unified under the framework of continuous-time generative modeling.
Continuous Normalizing Flows (CNFs) generalize discrete flows by formulating the transformation as the solution to an ordinary differential equation (ODE), yielding continuous-time invertible dynamics. Building on CNFs, Flow Matching techniques [71] learn velocity fields that directly transport samples from the base to the target distribution by minimizing an optimal-transport–informed loss, effectively bridging the gap between likelihood-based flows and transport-based formulations. In the realm of Text-to-Speech (TTS), models like VoiceBox [64] combine flow-matching objectives with scalable in-context learning to refine alignment trajectories. Similarly, Matcha-TTS [78] employs conditional flow matching guided by optimal transport to enhance style and prosody control.
In parallel, Diffusion Models utilize stochastic differential equations (SDEs) to gradually denoise samples, transforming them from pure noise into coherent data. A notable example in TTS is Grad-TTS [87], which leverages score-based diffusion in the spectrogram domain to achieve high-fidelity voice generation through iterative refinement.
Crucially, Flow Matching provides a principled link between deterministic Ordinary Differential Equation (ODE) flows and the stochastic SDE approaches characteristic of diffusion. By matching marginal distributions along continuous trajectories, flow matching unifies these two methodologies, combining the potential for exact density evaluation with the strong sample quality often associated with diffusion-based generative modeling.
Transformer-Based Model: The Transformer architecture [123] has revolutionized the field of text-to-speech, largely replacing earlier recurrent or convolutional approaches. Its core strength lies in the powerful self-attention mechanism within its encoder-decoder structure, enabling the effective conversion of input text sequences into compressed intermediate representations. These representations are then decoded to generate acoustic features, such as mel-spectrograms, which form the basis for fake voice generation [116]. This paradigm shift reveals in the development of numerous influential models. An early key example is Transformer-TTS [69], which successfully adapted the original Transformer by integrating it with design principles from the established Tacotron 2 architecture [102]. Seeking to overcome the latency issues inherent in autoregressive generation, the FastSpeech series emerged as a significant advancement. FastSpeech [97] pioneered non-autoregressive parallel generation for faster synthesis, relying on an external aligner to predict phoneme durations. Its successor, FastSpeech 2 [96], further refined this parallel approach by directly predicting variance information like pitch and energy within the model, simplifying the training pipeline and enhancing controllability. Subsequent research focused on improving the expressiveness and naturalness of the generated speech; DelightfulTTS [74], for instance, incorporated Conformer modules to better capture both local and global acoustic context. More recently, architectures like E2-TTS [27] explore novel directions such as flow-matching transformers operating directly on simple character sequences augmented with filler tokens, demonstrating the ongoing evolution of the Transformer acoustic model in TTS fake voice generation.
2.1.2 Vocoder
Vocoders are used for generating waveforms from acoustic features. We investigate vocoders based on two main categories: autoregressive vocoder and non-autoregressive vocoder. For latter category, we focus on three principal categories: Flow-Based, Diffusion-Based, and Generative Adversarial Network (GAN)-Based vocoders.
Autoregressive Vocoder: Unlike conventional parametric vocoders dependent on signal processing methods, autoregressive vocoders create audio samples sequentially, each new sample conditioned on prior produced samples and acoustic properties. By capturing the temporal dependencies in speech signals, Autoregressive vocoders such as WaveNet [122], WaveRNN [48], and LPCNet [121] have demonstrated remarkable capabilities in reproducing natural speech characteristics, including phase information and fine spectral details that are often lost in conventional vocoding methods. Their capacity to replicate the probabilistic character of voice waveforms has made them as foundation for contemporary neural vocoder.
Flow-Based Vocoder: Normalizing flow-based generative vocoders simplify complex probability distributions using invertible transformations, typically based on Gaussian distributions [116]. The core mechanism of these models converts data between complex and standard distributions via invertible mappings, sampling from the standard distribution and applying inverse transformations for generation. Foundation models in this approach include WaveGlow [88], which combined Glow and WaveNet for a non-autoregressive model using affine coupling layers, and WaveFlow [86], which leverages invertible transformations to enable efficient and parallel generation.
Diffusion-Based Vocoder: Diffusion vocoders leverage denoising diffusion probabilistic models [39, 116, 147] to map data and latent distributions via forward and reverse processes. The core mechanism consists of a forward process that adds noise to clean audio via a Markov chain with Gaussian transitions, and a reverse process that recovers clean audio by estimating noise with a simplified loss function. Key models in this field include foundation works such as DiffWave [58], which uses versatile Markov chain generation as the keystone diffusion-based vocoder, and WaveGrad [13], which employs gradient-based noise estimation. Noise distribution improvements have been achieved with PriorGrad [66], which introduces adaptive data-dependent priors. Several acceleration techniques have emerged, including FastDiff [41], which develops dynamic noise scheduling, and BDDM [62], which utilizes bilateral denoising diffusion modeling. Advanced optimizations continue to refine these approaches, such as FreGrad [82], which employs wavelet-based sub-band decomposition for improved efficiency. Rectified flow also has been used in diffusion models such as RFWave [72], which adapts Rectified Flow with multi-band processing to address latency issues, further improving the efficiency and quality of fake voice generation.
GAN-Based Vocoder: GAN vocoders employ adversarial training with a discriminator to assess authenticity [67, 116], enabling high-quality fake voice generation. Foundation models include MelGAN [59], which uses multi-scale discriminators and hinge loss, and ParallelWaveGAN [139], which implements multi-resolution STFT losses. Architectural improvements followed with HiFi-GAN [57] introducing multiperiod discriminators and multireceptive field fusion, UnivNet [42] featuring multiresolution discriminators for full-band signals, and BigVGAN [67] innovating with periodic activations. Artifact reduction was addressed by FA-GAN [104] with anti-aliased deconvolution and Avocado [7] mitigating spectral leakage. Digital Signal Processing integration advanced through DSP-GAN [108], SiFi-GAN [142] implementing source-filter theory, and Vocos [106] with Fourier-based time-frequency representation.
@@ 2.2 Fully End-to-End Models
Fully end-to-end text-to-speech models generate waveforms directly from text, bypassing the traditional acoustic model and vocoder pipeline with intermediate spectrograms. This approach mitigates error propagation in cascaded structures and simplifies the training process while potentially improving overall speech quality [116]. By optimizing all components in a single framework, these models can better capture the complex relationships between text and speech, resulting in higher fidelity generations. Key models in this category include the VITS series, with VITS [54] serving as a keystone model based on conditional variational autoencoder for end-to-end waveform generation, and MB-ISTFT-VITS [49] introducing multi-band generation and ISTFT decoding for enhanced speech detail. Other notable contributions include Your-TTS [10], the first work proposing a multilingual approach in the zero-shot voice conversion within multi-speaker text-to-speech scope, NaturalSpeech [115], which utilizes a differentiable durator and memory-enhanced bidirectional modeling, and Mega-TTS [43], which uses VQ-GAN with disentangled modeling for content and prosody control.
@@ 2.3 Neural Audio Codec Language Models
Neural audio codecs utilize neural networks to compress audio waveforms into compact representations via an encoder and reconstruct them using a decoder [103]. These codecs are integral components of modern text-to-speech pipelines, balancing low bit rates for compression with high audio fidelity.
Typically, these codecs generate compact representations using discrete tokens. Prominent examples like SoundStream [145], Encodec [23], and Descript Audio Codec [60] employ Vector Quantized Variational AutoEncoder (VQ-VAE) architectures often combined with residual vector quantizers for efficient encoding.
In text-to-speech, a language model predicts sequences of these discrete audio tokens based on input text, potentially conditioned on speaker embeddings or audio prompts. The codec’s decoder then generates the final audio waveform from these predicted tokens [15]. Key text-to-speech generators with zero-shot voice conversion leveraging this paradigm include VALL-E [15] which employs a two-stage approach with autoregressive and non-autoregressive transformer models as a keystone text-to-speech fake voice generator of neural codec based language model which uses Encodec [23] as tokenizer. NaturalSpeech series (2-3) [103, 44] utilizes continuous vector representations to shorten token sequences and incorporates a factorized codec design.
Recent advancements further enhance capabilities and address specific challenges and strengthen zero-shot voice conversion capabilities:
  • XTTS [9] and CosyVoice [26]: Focuses on massively multilingual support and sophisticated voice modification features.
  • MaskGCT-TTS [130]: Introduces a mask-and-predict training strategy, strengthening zero-shot voice conversion and computational efficiency.
  • CLaM-TTS [55]: Leverages probabilistic Residual Vector Quantization for superior compression in terms of token sequence length.
  • VoiceCraft [83]: Proposes a novel token rearrangement technique to improve synthesis quality.
  • FireRedTTS series [35, 36]: Integrates a semantic-aware speech tokenizer for generation from text/audio prompts and introduces a streamable codec architecture using causal convolutions for real-time applications.
The high fidelity achieved by models like NaturalSpeech3 [44] raises significant concerns about voice identity protection and potential misuse for impersonation. This necessitates research into robust detection methods and watermarking techniques for fake voice generation. Balancing rapid innovation with responsible deployment is crucial for maintaining public trust and realizing the benefits of advanced fake voice generation technologies.

### v2

@@ Abstract
The rapid advancement of fake voice generation technology has ignited a race with detection systems, creating an urgent need to secure the audio ecosystem. However, existing benchmarks suffer from a critical limitation: they typically aggregate diverse fake voice samples into a single dataset for evaluation. This practice masks method-specific artifacts and obscures the varying performance of detectors against different generation paradigms, preventing a nuanced understanding of their true vulnerabilities. To address this gap, we introduce the first ecosystem-level benchmark that systematically evaluates the interplay between 17 state-of-the-art fake voice generators and 8 leading detectors through a novel one-to-one evaluation protocol. This fine-grained analysis exposes previously hidden vulnerabilities and sensitivities that are missed by traditional aggregated testing. We also propose unified scoring systems to quantify both the evasiveness of generators and the robustness of detectors, enabling fair and direct comparisons. Our extensive cross-domain evaluation reveals that modern generators, particularly those based on neural audio codecs and flow matching, consistently evade top-tier detectors. We found that no single detector is universally robust; their effectiveness varies dramatically depending on the generator’s architecture, highlighting a significant generalization gap in current defenses. This work provides a more realistic assessment of the threat landscape and offers actionable insights for building the next generation of detection systems.
@@ 1 Introduction
The evolution of fake voice generation has transformed digital communication, with applications from Text-to-Speech (TTS) and voice conversion to audio reconstruction achieving unprecedented naturalness and quality [1]. Early techniques based on parametric [2, 3, 4, 5] or concatenative [6, 7, 8] methods have given way to deep neural architectures. Models like WaveNet [9] and Tacotron [10], and more recently, vocoders based on generative adversarial networks (GANs) [11] and diffusion models [12, 13, 14], have pushed the boundaries, achieving near-human levels of prosody and intelligibility.
This rapid advancement enables a wide range of beneficial services, from virtual assistants to content creation [15, 16]. However, as synthetic voices approach human parity, they also pose new and serious threats across various sectors. Beyond challenging the integrity of audio evidence in forensics, high-fidelity voice cloning is actively used for sheer deception. Malicious actors have targeted political figures [17] and facilitated large-scale financial fraud by impersonating executives [18]. These incidents highlight critical vulnerabilities in systems that rely on voice for authentication and verification.
@@
Consequently, an intense race has emerged between advanced fake voice generation techniques and the automated detection systems designed to counter them [19]. These detectors are critical Counter-Measures (CM) for technologies like Automatic Speaker Verification (ASV), which are vulnerable to spoofing attacks [20, 21]. While prior studies have evaluated detection systems [22, 12, 23, 24, 25], they have a significant limitation: they conventionally aggregate fake voice samples into a single, homogeneous dataset for training and evaluation [26, 27, 28, 29, 30, 31, 32, 33, 34]. This practice ignores critical nuances, preventing the identification of unique acoustic artifacts inherent to each generation paradigm, such as those from different voice conversion or synthesis models, and masking variations in detector performance across generator types. Recent work has shown that acoustic models and vocoders leave distinct “fingerprints,” which can be used for source attribution [35], a critical capability for forensics [36], copyright protection [37], and developing general defense strategies [38].
@@
We introduce an ecosystem-level benchmark that systematically evaluates the interplay between modern fake voice generation attack and detection defense. Our framework challenges 17 State-Of-The-Art (SOTA) fake voice generator, spanning TTS, TTS with voice conversion, and audio reconstruction, with 8 leading fake voice detectors using a novel one-to-one evaluation protocol. Unlike conventional assessments, our approach evaluates each generator-detector pair individually to expose unique acoustic artifacts, method-specific sensitivities, and previously hidden vulnerabilities. We proposed new unified metric for generator and detector to generally compare the performance of both in diverse dimensions, and we also clarify the underlying reasons for detector performance discrepancies, enabling the community to pinpoint techniques that threaten defenses. Figure 1 outlines our evaluation pipeline.
@@
    Comprehensive Security Taxonomy: We provide a detailed taxonomy of modern fake voice generation and detection systems, framed through a security lens to identify critical architectural vulnerabilities.
@@
    Large-Scale Empirical Assessment: We conduct the most extensive cross-domain security evaluation to date, benchmarking 8 State-Of-The-Art detectors against 17 diverse fake voice generators. This one-to-one analysis reveals previously hidden, method-specific vulnerabilities that are obscured by traditional aggregated testing.
@@
    Unified Performance Metrics: We propose a novel and unified scoring system to standardize the evaluation of both generator evasiveness and detector robustness, enabling fair comparisons across different technologies.
@@
    Actionable Recommendations: Based on our extensive empirical findings, we offer concrete recommendations for developing the next generation of resilient detection systems and designing security-aware voice generators.
@@ 2 Fake Voice Generators
This section provides a taxonomy of fake voice generation systems, focusing on Text-to-Speech (TTS), the process of converting text into spoken voice. We particularly examine zero-shot voice conversion, a technique enabling voice generation from a short voice clip of a target speaker without specific prior training [39].
The evolution of these systems is depicted in Figure 2. Fake voice generation has progressed from traditional two-stage pipelines to more integrated models. Traditional approaches use an acoustic model to convert text into an intermediate acoustic representation (like a spectrogram), which a vocoder then synthesizes into a waveform. These can be optimized as a single, fully end-to-end (E2E) model. A more recent paradigm involves neural audio codecs, which use a data-driven approach to compress and reconstruct voice. In this setup, a language model predicts a sequence of discrete voice tokens from text, which the codec’s decoder then converts into the final waveform.
@@
Figure 2: The evolution of fake voice generation. Traditional neural TTS pipelines convert text to acoustic representations, then use vocoders to synthesize waveforms. Fully E2E models optimize all components at once. Neural codec models learn to encode and decode audio, replacing static compression algorithms. Audio inputs are often used for speaker conditioning in zero-shot TTS.
2.1 Traditional Generators with Vocoders
Traditional fake voice generators consist of two main components: an acoustic model and a vocoder.
2.1.1 Acoustic Models
Acoustic models generate acoustic representations from linguistic features [23]. Key architectures include:
Autoregressive Models: These models, pioneered by the Tacotron series [10], generate spectrograms sequentially using encoder-attention-decoder architectures. While known for their naturalness, they face challenges with latency. Later models like DurIAN [40] improved alignment stability by using duration predictors instead of attention mechanisms.
Flow-Matching & Diffusion-Based Models: These models transform a simple distribution (e.g., Gaussian noise) into a complex data distribution. Flow-matching models learn a vector field to transport samples from a base to a target distribution, as seen in Matcha-TTS [41], which enhances prosody control. Diffusion models, such as Grad-TTS [42], gradually denoise samples from pure noise into coherent data, achieving high-fidelity synthesis through iterative refinement. Flow matching effectively unifies these deterministic (ODE-based) and stochastic (SDE-based) approaches [43].
Transformer-Based Models: The Transformer architecture [44] has become dominant in TTS due to its self-attention mechanism. Early examples like Transformer-TTS [45] adapted the architecture for voice. To address the latency of autoregressive methods, the non-autoregressive FastSpeech series [46, 47] introduced parallel generation by predicting phoneme durations, pitch, and energy, which simplified training and improved control.
2.1.2 Vocoders
Vocoders synthesize raw voice waveforms from the acoustic features generated by the acoustic model.
Autoregressive Vocoders: Models like WaveNet [9] and WaveRNN [48] generate voice samples sequentially, where each sample is conditioned on previous ones. This approach captures temporal dependencies well, resulting in highly natural voice but at the cost of slow inference speed.
Non-Autoregressive Vocoders: To accelerate generation, several parallel vocoding methods have been developed:
  • Flow-Based Vocoders, such as WaveGlow [49], use invertible transformations to map a simple distribution to a complex waveform, enabling fast, parallel synthesis.
  • Diffusion-Based Vocoders, like DiffWave [50] and the more efficient RFWave [51], generate voice by reversing a diffusion process that gradually adds noise. They offer high-quality synthesis with improving speeds.
  • GAN-Based Vocoders employ adversarial training. Models like MelGAN [14] and the highly influential HiFi-GAN [13] use a generator and one or more discriminators to achieve a balance of high fidelity and computational efficiency widely used in modern TTS systems.
@@ 2.2 Fully End-to-End Models
Fully end-to-end models generate waveforms directly from text, bypassing intermediate acoustic representations like spectrograms. This integrated approach simplifies training, avoids error propagation between stages, and can improve overall voice quality [23]. Key models such as VITS [52] use a conditional variational autoencoder with adversarial training to achieve high-fidelity E2E generation. Other notable contributions include Your-TTS [53], which pioneered multilingual zero-shot voice conversion, and NaturalSpeech [54], which introduced novel components like a differentiable durator to enhance naturalness.
@@ 2.3 Neural Audio Codec Language Models
A leading paradigm in modern TTS involves neural audio codecs and language models. A neural audio codec (e.g., Encodec [55]) uses an encoder to compress waveforms into a compact sequence of discrete tokens and a decoder to reconstruct the voice from these tokens. A language model is then trained to predict these voice tokens from input text, often conditioned on a short voice for zero-shot voice conversion.
This approach was popularized by VALL-E [39], which uses a Transformer-based language model to predict Encodec tokens. Subsequent models have advanced the generator performance:
  • NaturalSpeech 2 & 3 [56, 57]: Introduced a factorized codec design and continuous vector representations to achieve extremely high voice fidelity.
  • XTTS [58] & CosyVoice [59]: Focused on massively multilingual support and fine-grained voice cloning capabilities.
  • VoiceCraft [60]: Proposed a novel token rearrangement technique to improve synthesis quality and editing capabilities.
  • FireRedTTS [61]: Developed a streamable architecture with causal convolutions, enabling real-time applications.
The high fidelity achieved by these models, particularly their ability to convincingly clone voices from just a few seconds of voice, raises significant concerns about misuse and impersonation. This underscores the urgent need for robust fake voice detection and watermarking techniques to ensure the responsible deployment of these powerful technologies.